OpenStudy (anonymous):

The diameter of a circle has endpoints whose coordinates are R(-2, 2) and S(4, 2). Find the equation of the circle. (x - )^2 + (y - )^2

OpenStudy (anonymous):

(x-4)(x+2) + (y-2)(y-2) = 0 this is the equation of the circle, you only need expand it now.

OpenStudy (campbell_st):

find the mid point.... this will be the centre (h, k).... find the distance between the two points R and S and then half it.... this is the radius (r)... Sub into \[(x - h)^2 + (y -k)^2 = r^2\]